BioCom, Inc.: Part​ 3, A Fresh Look at the WACC

In the course of discussing the​ fiber-optic blood pressure monitor project that we introduced in Chapter​ 10, a recently hired financial analyst who is working on her MBA asks how the company arrived at 9.5 % as the discount rate to use when evaluating capital budgeting projects. Her question is followed by an embarrassing silence that seems to last forever. ​ Eventually, the​ comptroller, who has been with the company for many​ years, offers an explanation. When the company first began to use discounted cash flow methods for capital budgeting decisions in the​ 1980s, it hired a consultant to explain internal rate of return and net present value. The consultant used 10.5 % in all his​ examples, so BioCom did the same. By the late​ 1990s, interest rates had fallen​ considerably, and the company was rejecting some seemingly profitable projects because the hurdle rate was too​ high, so it lowered it to 9.5 % As far as he​ knew, that was the end of the story.

Interestingly, many participants in the discussion—even those who are not from accounting and finance—are aware of the weighted average cost of capital and have a good idea of how to compute​ it, but no one had ever attempted to do so for BioCom. After a brief​ discussion, they ask the person who raised the question in the first place to analyze the​ company's debt and equity and report back in a week with her estimate of the​ company's weighted average cost of capital.  

She begins by gathering the following​ information: BioCom has two outstanding bond issues. Bond 1 matures in six​ years, has a par value of ​$1,000​, has a coupon rate of

7.6 % paid​ semiannually, and now sells for $1,031.85. Bond 2 matures in sixteen​ years, has a par value of $1,000​, has a coupon rate of 8.4% paid​ semiannually, and now sells for $1,034.25.

The preferred stock has a par value of ​$50​, pays a dividend of $1.521.52​, and has a current market value of $19.15. The common stock sells for $35.45 per share and recently paid a dividend of ​$2.552.55. The company expects dividends to grow at an average annual rate of 6.2% for the foreseeable future. The​ risk-free rate is 3.2%​, the expected rate of return on the market portfolio is 12.3%​, ​BioCom's beta is 1.25​, and its marginal tax rate is 34%.

​ BioCom's capital structure is shown in the popup​ window:

  Type of Capital

Percent of Book Value

Percent of Market Value

  Bond 1

20​%

17​%

  Bond 2

22​%

19​%

  Preferred stock

21​%

12​%

  Common stock

7​%

Not available

  Retained earnings

30​%

Not available

  Total common equity

52​%

                                               

Assist the financial analyst by answering these questions.

Questions

1.  Compute the yield to maturity and the​ after-tax cost of debt for the two bond issues.

2.  Compute​ BioCom's cost of preferred stock.

3.  Compute​ BioCom's cost of common equity. Use the average of results from the dividend growth model and the security market line.

4.  Compute​ BioCom's weighted average cost of capital. Should you use book values or market values for this​ computation?

5.  BioCom could sell new bonds with maturities of twenty years at approximately the same yield as bond 2. It​ would, however, incur flotation costs of

​$15 per $1,000 of par value. Estimate the effective interest rate BioCom would have to pay on a new issue of​ long-term debt.

6.  Some of​ BioCom's projects are low​ risk, some average​ risk, and some high risk. Should BioCom use the same cost of capital to evaluate all its​ projects, or should it adjust the discount rate to reflect different levels of​ risk?
Answer #1

1.Sol)

Yield to maturity of both bonds can be calculated in excel as

BS File Home Insert Page Layout Formulas Data Review View Help Search pyn - ab Wrap Text Merge & Center - Paste & Cut Calibri

After tax cost debt of Bond 1 = 6.94(1-0.34) = 4.58%

After tax cost of Bond 2 = 8.02*(1-0.34) = 5.29%

2.Sol)

Dividend on preferred stock = 1.52

Price of preferred stock = 19.15

Cost of preferred stock = Dividend / price

= 1.52/19.15 = 7.93%

3.Sol)

Cost of equity using dividend growth model can be calculated from the following formula

Price of stock = Dividend / (Cost of equity - Growth rate)

35.45 = 2.55 / (Cost of equity - 0.062)

Cost of equity = 2.55/35.45 + 0.062

Cost of equity = 13.39%

Cost of equity using Security market line

Cost of equity = Rf + Beta*(Rm-Rf)

= 3.2 + 1.25*(12.3-3.2)

= 14.575%

Average of cost of common equity resulted from both methods = (13.39+14.575)/2 = 13.98%

4.Sol)

Market value of WACC should be preferred over Book value, since it helps in evaluating new projects more appropriately.

WACC = Cost of Bond1*Weight of Bond1 + Cost of Bond 2* Weight of Bond2 + Cost of preferred stock* Weight of Preferred stock + Cost of common equity* Weight of common equity

= 4.58*0.17 + 5.29*0.19 + 7.93*0.12 + 13.98*0.52

= 0.7786 + 1.0051 + 0.9516 + 7.2696

= 10%